OpenStudy (niahloves):

Translate the word phrase into a math expression. Antwan pays $12 per hour to rent a canoe for the afternoon. A. 12 + h B. 12 · h C. 12 ÷ h D. h ÷ 12

OpenStudy (anonymous):

it is A if you need help with anymore i'm here :)

OpenStudy (niahloves):

ok thanks

OpenStudy (niahloves):

@mathstudent55 what do you think?

OpenStudy (mathstudent55):

Let's figure it out with numbers. If the cost is $12 per hour, how much do you pay in 1 hour?

OpenStudy (niahloves):

:?

OpenStudy (niahloves):

;/

OpenStudy (mathstudent55):

It costs $12 per hour to rent a canoe. You want the canoe for exactly one hour. How much will it cost you for 1 hour?

OpenStudy (niahloves):

12?

OpenStudy (mathstudent55):

Correct. Now let's say you want to rent he canoe for 2 hours, not just 1 hour. How much will it cost for 2 hours? The cost is still $12 per hour, but you want 2 hours not 1. How much will the 2 hours cost?

OpenStudy (niahloves):

$24

OpenStudy (mathstudent55):

Great. So far we have: Renting for 1 hour: 12 * 1 = 12 Renting for 2 hours: 12 * 2 = 24 How about if you want to rent for 3 hours?

OpenStudy (niahloves):

Then it will be 12*3=36

OpenStudy (mathstudent55):

Exactly. Notice that in each case, we multiply the price per hour, $12, by the number of hours. If we want 5 hours, it's 12 * 5 For 10 hours it's 12 * 10 What if you want it for h hours?

OpenStudy (niahloves):

it will be 12*h

OpenStudy (mathstudent55):

\(\Huge \color{red}{B~I~N~G~O~!!!}\) You are correct. Which choice is 12*h?

OpenStudy (mathstudent55):

As you can see, the answer is not A.

OpenStudy (mathstudent55):

Ok, gtg. Bye.

OpenStudy (niahloves):

bye